Please can someone help me solve this problem?
In slope intercept form, give the equation of the line that contains
(1-2),(-2,1)

Thank you for your help

Answer by mananth(16946)

x1 y1 x2 y2
1 -2 -2 1

slope m = (y2-y1)/(x2-x1)
( 1 - -2 )/( -2 - 1 )
( 3 / -3 )
m= - 5/ 8

Plug value of the slope and point ( 1,-2) in
Y = m x + b
-2.00 = - 5/8 + b
b= -2 - - 5/8
b= -11/ 8
So the equation will be
Y = - 5/8 x -11/8
m.ananth@hotmail.ca